OpenStudy (jhannybean):

Calculus 3: Simplifying \[\int_{0}^{\pi/4} \int_{0}^{\pi/2} \int_{0}^{\sqrt{2}} (\rho\sin\phi \cos\theta)(\rho \sin \phi \sin \theta)\rho^2\sin\phi d\rho d\theta d\phi\]

OpenStudy (anonymous):

\[\int_{0}^{\pi/4} \int_{0}^{\pi/2} \int_{0}^{\sqrt{2}} (\rho\sin\phi \cos\theta)(\rho \sin \phi \sin \theta)\rho^2\sin\phi d\rho d\theta d\phi\] Reducing to what?

OpenStudy (dan815):

first multiply all

OpenStudy (jhannybean):

\[\int_{0}^{\pi/4} sin^3\phi d\phi \cdot \int_{0}^{\pi/2}\sin\theta\cos\theta \cdot \int_{0}^{\sqrt{2}} \rho^4 d\rho\]

OpenStudy (jhannybean):

crap this takes a while to write out.

OpenStudy (dan815):

oh its separte integrals like that what?

OpenStudy (jhannybean):

I'm not really stuck on the process of simplifying it, just the identities that come along with some of the integration, I have forgotten :(

OpenStudy (dan815):

whats your actual question in latex form

OpenStudy (dan815):

okay so

OpenStudy (jhannybean):

I just ocnfused @radar and @eliassaab , sorry, hahaha

OpenStudy (dan815):

p^4*coss...sin... root2^5/5*co..sin,, ---------------- and u got sin phi that factors nicely outta the cos and sin theta

OpenStudy (dan815):

so u get some factor *costheta+sintheta

OpenStudy (dan815):

wait no i mean cos theta* sin theta

OpenStudy (dan815):

and cos theta sin theta can be rewritten as

OpenStudy (dan815):

as the sum of addition and difference identities

OpenStudy (jhannybean):

\[\int_{0}^{\pi/4} sin^3\phi d\phi \cdot \int_{0}^{\pi/2}\sin\theta\cos\theta \cdot \int_{0}^{\sqrt{2}} \rho^4 d\rho\]\[\int_{0}^{\pi/4} (1-cos^2\phi)\sin\phi d\phi \cdot \int_{0}^{\pi/2} \color{red}{sin\theta\cos\theta} \cdot \left.\frac{1}{5}\rho^5\right]_{0}^{\sqrt{2}}\]

OpenStudy (dan815):

sin(a+b) + sin(a-b) --------------- 2

OpenStudy (dan815):

or something like that lets see

OpenStudy (dan815):

ya umm u know how sin theta and cos theta pop out..

OpenStudy (jhannybean):

That red portion... would it be....

OpenStudy (dan815):

sin(A + B) = sin(A)cos(B) + cos(A)sin(B) sin(A-b)= sin(a)cos(b) - cos(a)sin(b) sin(a+b)+sin(a-b)= 2 sin(a)cos(b)

OpenStudy (jhannybean):

Well i know \(sin(2\theta) = 2sin(\theta)cos(\theta)\)

OpenStudy (dan815):

so we can simplify sin(a)cos(b)

OpenStudy (dan815):

with that process

OpenStudy (dan815):

right a-b =0 xD

OpenStudy (dan815):

when a=b

OpenStudy (dan815):

sin(a+b)+sin(a-b)= 2 sin(a)cos(b) sin(a+a)+sin(a-a)=2sin(a)cos(a) sin(2a) ------ = sin a cos a 2

OpenStudy (dan815):

okay now just integrate

OpenStudy (jhannybean):

Oh wait...yeah i was just about to type that. haha

OpenStudy (zarkon):

just use a u-sub

OpenStudy (jhannybean):

For which portion?

OpenStudy (dan815):

hes saying u can use u sub for cos theta sin theta

OpenStudy (zarkon):

\[\sin(\theta)\cos(\theta)\]

OpenStudy (jhannybean):

Oh. Hmm. I was thinking I'd use a u-sub for the more complicated one, the first integral

OpenStudy (dan815):

u=sintheta du=costheta dtheta so sin a cos a da = u du

OpenStudy (dan815):

simplifies nicely

OpenStudy (dan815):

but u already know the identity now so

OpenStudy (anonymous):

To check your final answer \[ \int_0^{\frac{\pi }{4}} \left(\int_0^{\frac{\pi }{2}} \left(\int_0^{\sqrt{2}} \rho \rho \rho ^2 \sin (\theta ) \cos (\theta ) \sin (\phi ) \sin (\phi ) \sin (\phi ) \, d\rho \right) \, d\theta \right) \, d\phi =\frac{1}{15} \left(4 \sqrt{2}-5\right) \]

OpenStudy (anonymous):

\[\int_0^{\frac{\pi }{4}} \left(\int_0^{\frac{\pi }{2}} \left(\int_0^{\sqrt{2}} \rho \rho \rho ^2 \sin (\theta ) \cos (\theta ) \sin (\phi ) \sin (\phi ) \sin (\phi ) \, d\rho \right) \, d\theta \right) \, d\phi =\\\frac{1}{15} \left(4 \sqrt{2}-5\right) \]

OpenStudy (jhannybean):

\[\int_{0}^{\pi/4} (1-cos^2\phi)\sin\phi d\phi \cdot \int_{0}^{\pi/2} \color{red}{sin\theta\cos\theta} \cdot \left.\frac{1}{5}\rho^5\right]_{0}^{\sqrt{2}}\]\[\int_{0}^{\pi/4} (1-cos^2\phi)sin \phi d\phi\] If we let u = cos\(\theta\), du = -sin\(\theta\)d\(\theta\)\[-\int_{0}^{\pi/4} (1-u^2)du\] The first part. right?

OpenStudy (dan815):

so final problem is intergral of sinphi ^ 3

OpenStudy (dan815):

yah thats rght

OpenStudy (zarkon):

no

OpenStudy (zarkon):

change your limits of integration after substitution

OpenStudy (jhannybean):

Do you have to.... or can you resub before you integrate the limits?

OpenStudy (zarkon):

the way you have it written is incorrect

OpenStudy (zarkon):

if you want you could write \[-\int_{u(0)}^{u(\pi/4)} (1-u^2)du\]

OpenStudy (dan815):

\[\int_{0}^{\pi/4} (1-cos^2\phi)\sin\phi d\phi \cdot \int_{0}^{\pi/2} \color{red}{sin\theta\cos\theta} \cdot \left.\frac{1}{5}\rho^5\right]_{0}^{\sqrt{2}} d\theta \]

OpenStudy (jhannybean):

\[-\int_{1}^{\sqrt{2}/2} (1-u^2)du\]

OpenStudy (dan815):

int cos^2a* sin a da cos^3a = ----- 3

OpenStudy (jhannybean):

Ahh.. I see

OpenStudy (dan815):

cos(pi/4)=1/root2

OpenStudy (anonymous):

To Check your answer \[\int_0^{\frac{\pi }{4}} \sin ^3(\phi ) \, d\phi =\frac{1}{12} \left(8-5 \sqrt{2}\right) \\ \int_0^{\frac{\pi }{2}} \sin (\theta ) \cos (\theta ) \, d\theta =\frac{1}{2}\\ \int_0^{\sqrt{2}} \rho ^4 \, d\rho =\frac{4 \sqrt{2}}{5}\]

OpenStudy (jhannybean):

1. \[\left. -u+\frac{1}{3}u^3\right]_{1}^{\sqrt{2}/2}\]2. u= sin\(\theta\), du = \(\cos\theta d\theta\)\[\int_{0}^{1}udu = \left.\frac{1}{2}u^2\right]_{0}^{1}\]3.\[\left.\frac{1}{5}\rho^5\right]_{0}^{\sqrt{2}}\]Haven't simplified them but this is what I have with both of the u-subs.

OpenStudy (jhannybean):

Oh cool, it matches up with yours. @eliassaab

OpenStudy (jhannybean):

Solving it I'd get 1. \[\left(-\frac{\sqrt{2}}{2} +\frac{1}{3}\left(\frac{\sqrt{2}}{2}\right)^3\right)-\left(1+\frac{1}{3}\right)\]

OpenStudy (jhannybean):

But yeah I get it.... thank you guys!!

OpenStudy (jhannybean):

and theres a typo there, meant -1*